OpenStudy (shamallamadingdong):

Help Which is the equation of the line that contains points (8, 10) and (–4, 2)?

OpenStudy (aaronandyson):

Find the slope.

OpenStudy (shamallamadingdong):

y-8=2/3(x-10) y-10=2/3(x-8) y-10=3/2(x-8) y-8=3/2(x-10)

OpenStudy (shamallamadingdong):

slope wouldent that be 12 over 6

OpenStudy (aaronandyson):

\[m = y_2 - y_1/x_2-x_1\]

OpenStudy (aaronandyson):

m = (y2-y1)/(x2-x1) m = (10-2)/[8-(-4)] m = 8/12 m = ?

OpenStudy (aaronandyson):

Note: - of - is +

OpenStudy (shamallamadingdong):

what i am so confuzzled rn

OpenStudy (shamallamadingdong):

i understand that -of-=+

OpenStudy (shamallamadingdong):

but in the end wouldent the slope be 8/12?

OpenStudy (aaronandyson):

8/12 simplifies to 2/3

OpenStudy (shamallamadingdong):

so now i am down to two answers....... wouldent it be B?

OpenStudy (aaronandyson):

CORRECT!

OpenStudy (shamallamadingdong):

YAAASSSSS